OpenStudy (anonymous):

If you were to use the substitution method to solve the following system, choose the new equation after the expression equivalent to y from the second equation is substituted into the first equation. 2x - 5y = -3 6x - 3y = -3 (4 points) Question 8 answers 2(2x + 1) - 5y = -3 2(-2x - 1) - 5y = -3 2x - 5(2x + 1) = -3 2x - 5(-2x - 1) = -3

OpenStudy (anonymous):

First use the second equation and solve for y. 6x-3y=-3 -3y=-3-6x \[y=\frac{ -3-6x }{ -3 }\] y=1+2x or y=2x+1 Now insert your new answer for y in the first equation. 2x-5(2x+1)=-3

OpenStudy (anonymous):

Does that make sense?